# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ A.2 Cubic action from scalar-tensor-tensor interactions
$\quad\quad\quad\quad\quad\quad\ -\frac{\mu}{2a^{2}}\alpha\dot{h}_{ij}\partial^{2}h_{ij}-\frac{\mathcal{G}_{T}}{4a^{2}}\partial_{k}\beta\dot{h}_{ij}\partial_{k}h_{ij}-\frac{\mu}{2a^{2}}\biggl{(}\partial_{i}\partial_{j}\beta\dot{h}_{ik}\dot{h}_{jk}-\frac{1}{2}\partial^{2}\beta\dot{h}_{ij}^{2}\biggr{)}\biggr{]},$ (A.10)
@@
$S^{(3)}_{shh}=\int{\rm d}t{\rm d}^{3}x\biggl{[}a^{3}\mathcal{L}_{shh}+E_{shh}\biggr{]},$ (A.14)
where $\mathcal{L}_{shh}$ is given in Eq. (3.9), and $E_{shh}$ is defined by
$E_{shh}=\frac{\mu}{4\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}E^{s}+\frac{\mathcal{G}_{T}^{2}}{2\Theta\mathcal{F}_{T}}\biggl{(}\frac{\zeta}{2}+\frac{\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr{)}\dot{h}_{ij}E^{h}_{ij}.$ (A.15)
Each coefficient of the terms in $\mathcal{L}_{shh}$ takes the following form,
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl{\{}1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl{[}-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}\biggr{]}\biggr{\}},$ (A.16)
$b_{2}$ $=\frac{\mathcal{F}_{S}}{8},$ (A.17)
$b_{3}$ $=-\frac{\mathcal{G}_{S}}{4},$ (A.18)
$b_{4}$ $=\frac{\mathcal{G}_{T}}{8\Theta\mathcal{F}_{T}}(\mathcal{G}_{T}^{2}-\Gamma\mathcal{F}_{T})+\frac{\mu}{4}\biggl{[}\frac{\mathcal{G}_{S}}{\mathcal{G}_{T}}-1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}(6+g_{S})\biggr{]}+\frac{\mathcal{G}_{T}^{2}}{4}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mu}{\Theta\mathcal{F}_{T}}\biggr{)},$ (A.19)
$b_{5}$ $=\frac{\mu\mathcal{G}_{T}}{4\Theta}\biggl{(}\frac{\mathcal{F}_{S}\mathcal{G}_{T}}{\mathcal{F}_{T}\mathcal{G}_{S}}-1\biggr{)},$ (A.20)
$b_{6}$ $=-\frac{\mu\mathcal{G}_{S}}{2\mathcal{G}_{T}},$ (A.21)
$b_{7}$ $=\frac{\mu\mathcal{G}_{T}}{2\Theta}.$ (A.22)
@@
$h_{ij}$ $\to h_{ij}+\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\biggl{(}\zeta+\frac{2\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr{)}\dot{h}_{ij},$ (A.23)
$\zeta$ $\to\zeta+\frac{\mu}{8\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}.$ (A.24)
@@ Appendix B Primordial STT bispectrum
In this section, we summarize the details of the computations of the primordial STT bispectrum. To obtain a general expression in the context of the non-Bunch-Davies states, we assume that both scalar and tensor modes are in non-Bunch-Davies states. In the main text, we chose $\eta_{0}$ in such a way that the primordial GWs at the small scales avoid the Trans-Planckian problem. However, the results obtained under Eq. (3.1) are valid at least for the primordial non-Gaussianities on CMB scales. Thus, to obtain a general expression that can be used widely, we consider possible values of $\eta_{0}$. We first assume that (i) the short modes of the tensor perturbations were initially in excited states, i.e., $\eta_{0}$ denotes the time when both long and short modes are on subhorizon scales, $-k_{i}\eta_{0}\gg 1$ (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{i}^{\prime}\eta_{0}\gg 1$ under the squeezed limit). Then, we consider the case in which (ii) the perturbations got excited from the Bunch-Davies state when the long mode crossed the sound horizon, $-c_{s}k_{1}\eta_{0}=1$ (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}=1$ under the squeezed limit). We also consider the case where (iii) the long mode was already on superhorizon scales, $-k_{L}\eta_{0}\ll 1$, while the short modes were still on subhorizon scales, $-k_{S}\eta_{0}\gg 1$, when the perturbations got excited (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}\ll 1$ under the squeezed limit). Last, we assume that (iv) both long and short modes were on superhorizon scales, $-k_{i}\eta_{0}\ll 1$, when the perturbations got excited (i.e., $-K^{\prime}\eta_{0}\ll 1,-K_{i}^{\prime}\eta_{0}\ll 1$ under the squeezed limit).
@@
$\quad\quad\quad\quad\ +(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\Biggr{]}.$ (B.1)
@@ B.1 $b_{1}$ term
$\quad\quad\quad\quad\quad+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.2)
@@
$\mathcal{I}_{b_{1},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (B.4)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (B.5)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (B.6)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.7)
@@
In the case of (i), we obtain
$\mathcal{I}_{b_{1},K^{\prime}}$ $=\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}),$ (B.8)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(-k_{1},k_{2},k_{3}),$ (B.9)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (B.10)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (B.11)
@@
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}}.$ (B.12)
In the cases of (ii), (iii), and (iv), $\mathcal{I}_{b_{1},K^{\prime}}$ and $\mathcal{I}_{b_{1},K_{1}^{\prime}}$ take the same form with Eq. (B.8) and Eq. (B.9), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (B.13)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (B.14)
@@
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}).$ (B.15)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
@@ B.2 $b_{2}$ term
$\quad\quad\quad\ +\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.16)
@@
$\mathcal{V}_{b_{2}}=\frac{k_{1}^{2}-k_{2}^{2}-k_{3}^{2}}{2}\mathcal{V}_{b_{1}},$ (B.17)
@@
$\mathcal{I}_{b_{2},K^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (B.18)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(-i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (B.19)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(-i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (B.20)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(-i+c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta},$ (B.21)
In the case of (i), we obtain
$\mathcal{I}_{b_{2},K^{\prime}}$ $=\mathcal{J}_{b_{2}}(k_{1},k_{2},k_{3}),$ (B.22)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(-k_{1},k_{2},k_{3}),$ (B.23)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (B.24)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (B.25)
@@
$\quad\quad\quad\quad\ +c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})],$ (B.26)
and we used the fact that the term proportional to $1/\eta_{*}$ with $\eta_{*}(\to 0)$ denoting the conformal time at the end of inflation does not contribute to the three-point function because of the following equation,
@@
×$\quad\ \times\left(\alpha^{(s_{3})*}_{k_{3}}-\beta^{(s_{3})*}_{k_{3}}\right)\biggr{]}=0.$ (B.27)
In the other three cases, $\mathcal{I}_{b_{2},K^{\prime}}$ and $\mathcal{I}_{b_{2},K_{1}^{\prime}}$ take the same form with Eq. (B.22) and Eq. (B.23), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (B.28)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (B.29)
@@
$\quad\quad\quad\quad\ +\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr{]}.$ (B.30)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.

### v3

@@ A.2 Cubic action from STT interactions
$\quad\quad\quad\quad\quad\quad\ -\frac{\mu}{2a^{2}}\alpha\dot{h}_{ij}\partial^{2}h_{ij}-\frac{\mathcal{G}_{T}}{4a^{2}}\partial_{k}\beta\dot{h}_{ij}\partial_{k}h_{ij}-\frac{\mu}{2a^{2}}\biggl(\partial_{i}\partial_{j}\beta\dot{h}_{ik}\dot{h}_{jk}-\frac{1}{2}\partial^{2}\beta\dot{h}_{ij}^{2}\biggr)\biggr],$ (148)
@@
$S^{(3)}_{shh}=\int{\rm d}t{\rm d}^{3}x\biggl[a^{3}\mathcal{L}_{shh}+E_{shh}\biggr],$ (152)
where $\mathcal{L}_{shh}$ is given in Eq. (48), and $E_{shh}$ is defined by
$E_{shh}=\frac{\mu}{4\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}E^{s}+\frac{\mathcal{G}_{T}^{2}}{2\Theta\mathcal{F}_{T}}\biggl(\frac{\zeta}{2}+\frac{\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr)\dot{h}_{ij}E^{h}_{ij}.$ (153)
Each coefficient of the terms in $\mathcal{L}_{shh}$ takes the following form:
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl\{1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl[-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl(\frac{\mathcal{G}_{T}}{\Theta}\biggr)\biggr]\biggr\},$ (154)
$b_{2}$ $=\frac{\mathcal{F}_{S}}{8},$ (155)
$b_{3}$ $=-\frac{\mathcal{G}_{S}}{4},$ (156)
$b_{4}$ $=\frac{\mathcal{G}_{T}}{8\Theta\mathcal{F}_{T}}(\mathcal{G}_{T}^{2}-\Gamma\mathcal{F}_{T})+\frac{\mu}{4}\biggl[\frac{\mathcal{G}_{S}}{\mathcal{G}_{T}}-1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}(6+g_{S})\biggr]+\frac{\mathcal{G}_{T}^{2}}{4}\frac{{\rm d}}{{\rm d}t}\biggl(\frac{\mu}{\Theta\mathcal{F}_{T}}\biggr),$ (157)
$b_{5}$ $=\frac{\mu\mathcal{G}_{T}}{4\Theta}\biggl(\frac{\mathcal{F}_{S}\mathcal{G}_{T}}{\mathcal{F}_{T}\mathcal{G}_{S}}-1\biggr),$ (158)
$b_{6}$ $=-\frac{\mu\mathcal{G}_{S}}{2\mathcal{G}_{T}},$ (159)
$b_{7}$ $=\frac{\mu\mathcal{G}_{T}}{2\Theta}.$ (160)
@@
$h_{ij}$ $\to h_{ij}+\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\biggl(\zeta+\frac{2\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr)\dot{h}_{ij},$ (161)
$\zeta$ $\to\zeta+\frac{\mu}{8\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}.$ (162)
@@ Appendix B Primordial STT bispectrum
In this section, we summarize the details of the computations of the primordial STT bispectrum. To obtain a general expression in the context of the non-Bunch-Davies states, we assume that both scalar and tensor modes are in non-Bunch-Davies states. In the main text, we chose $\eta_{0}$ in such a way that the primordial GWs at the small scales avoid the Trans-Planckian problem. However, the results obtained under Eq. (36) are valid at least for the primordial non-Gaussianities on CMB scales. Thus, to obtain a general expression that can be used widely, we consider possible values of $\eta_{0}$. We first assume that (i) the short modes of the tensor perturbations were initially in excited states, i.e. $\eta_{0}$ denotes the time when both long and short modes are on subhorizon scales, $k_{i}\eta_{0}\gg 1$ (i.e. $K^{\prime}\eta_{0}\gg 1,-K_{i}^{\prime}\eta_{0}\gg 1$ under the squeezed limit). Then, we consider the case in which (ii) the perturbations got excited from the Bunch-Davies state when the long mode crossed the sound horizon, $c_{s}k_{1}\eta_{0}=1$ (i.e. $K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}=1$ under the squeezed limit). We also consider the case where (iii) the long mode was already on superhorizon scales, $k_{L}\eta_{0}\ll 1$, while the short modes were still on subhorizon scales, $k_{S}\eta_{0}\gg 1$, when the perturbations got excited (i.e. $K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}\ll 1$ under the squeezed limit). Last, we assume that (iv) both long and short modes were on superhorizon scales, $k_{i}\eta_{0}\ll 1$, when the perturbations got excited (i.e. $K^{\prime}\eta_{0}\ll 1,-K_{i}^{\prime}\eta_{0}\ll 1$ under the squeezed limit).
@@
$\quad\quad\quad\quad\ +(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\Biggr].$ (163)
@@ B.1 $b_{1}$ term
$\quad\quad\quad\quad\quad+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr$], (164)
@@
$\mathcal{I}_{b_{1},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (166)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (167)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (168)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (169)
@@
In case (i), we obtain
$\mathcal{I}_{b_{1},K^{\prime}}$ $=\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}),$ (170)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(-k_{1},k_{2},k_{3}),$ (171)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (172)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (173)
@@
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}}.$ (174)
In cases (ii), (iii), and (iv), $\mathcal{I}_{b_{1},K^{\prime}}$ and $\mathcal{I}_{b_{1},K_{1}^{\prime}}$ take the same form as in Eq. (170) and Eq. (171), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (175)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (176)
@@
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}).$ (177)
In case (ii), $\eta_{0}$ is replaced with $1/(c_{s}k_{1})$.
@@ B.2 $b_{2}$ term
$\quad\quad\quad\ +\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr$], (178)
@@
$\mathcal{V}_{b_{2}}=\frac{k_{1}^{2}-k_{2}^{2}-k_{3}^{2}}{2}\mathcal{V}_{b_{1}},$ (179)
@@
$\mathcal{I}_{b_{2},K^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (180)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(-i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (181)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(-i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (182)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(-i+c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta},$ (183)
In case (i), we obtain
$\mathcal{I}_{b_{2},K^{\prime}}$ $=\mathcal{J}_{b_{2}}(k_{1},k_{2},k_{3}),$ (184)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(-k_{1},k_{2},k_{3}),$ (185)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (186)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (187)
@@
$\quad\quad\quad\quad\ +c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})],$ (188)
and we used the fact that the term proportional to $1/\eta_{*}$, with $\eta_{*}(\to 0)$ denoting the conformal time at the end of inflation, does not contribute to the three-point function because of the following equation:
@@
$\quad\ \times\left(\alpha^{(s_{3})*}_{k_{3}}-\beta^{(s_{3})*}_{k_{3}}\right)\biggr]=0.$ (189)
In the other three cases, $\mathcal{I}_{b_{2},K^{\prime}}$ and $\mathcal{I}_{b_{2},K_{1}^{\prime}}$ take the same form as in Eq. (184) and Eq. (185), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (190)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (191)
@@
$\quad\quad\quad\quad\ +\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr].$ (192)
In case (ii), $\eta_{0}$ is replaced with $1/(c_{s}k_{1})$.